Need help please with thus equation : solve for z : z+i^3 = 1/1-i^2
Answer by rothauserc(4718)

i^2 = -1 and i^3 = -i
z +i^3 = 1/1-i^2
z -i = 1/2
2z -2i = 1
2z = 1 +2i
z = (1+2i) / 2